may you please help me figure out why this Q=3N+5, is the rule to this expression here. When you figure out this problem, may you please tell me how i can write a rule for this? Thank you vey much
N Q
2 11
4 17
6 23
8 29
10 35

Answer by nycsharkman(136)

may you please help me figure out why this Q = 3N + 5, is the rule to this expression here. When you figure out this problem, may you please tell me how I can write a rule for this? Thank you vey much
N Q
2 11
4 17
6 23
8 29
10 35
Let's look at the first-two values of N and Q.
What is the value of Q when N is replaced by 2 in the given formula
Q = 3N + 5? If N = 2, then Q = 11.
LOOK:
Q = 3(2) + 5
Q = 6 + 5
Q = 11
What is the value of Q when N = 4 in the given equation?
If N = 4, then Q = 17.
LOOK:
Q = 3(4) + 5
Q = 12 + 5
Q = 17
The same process applies to all values of N given in the N and Q list above.
The rule is the equation given: Q = 3N + 5. This equation becomes the key needed to find all the value for Q given in the list.